OpenStudy (anonymous):

QUESTION FOR FUN SOLVE FOR n: n^n +n^(n-1) + n^(n-2)+...+n^2+n^1+n^0 = n^n^n

OpenStudy (shubhamsrg):

1 + n + n^2 .. n^n = n^n^n n=1 is a clear solution for n not equal to 1, we have => (n^(n+1) -1)/(n-1) = n^n^n unsure what will come next.

OpenStudy (anonymous):

1^1 +1^0 = 1^1^1 2=1 which is FALSE

OpenStudy (anonymous):

So, n=1 is not a solution

OpenStudy (anonymous):

Should I post my solution?

OpenStudy (shubhamsrg):

Oh,,Sorry about that.

OpenStudy (shubhamsrg):

Well I failed to do it doesn't mean no one will do it. Many geniuses are there on OS. I'd recommend you to wait. :)

OpenStudy (anonymous):

As u say

OpenStudy (anonymous):

2 minutes, I'm having a crack.

OpenStudy (anonymous):

U can take your time

OpenStudy (anonymous):

Before I spend the next hour looking at it to find I cant find a solution, is there definately solutions?

OpenStudy (anonymous):

Apologise if it is obvious to have solutions, its been a long day of uni

OpenStudy (anonymous):

Actually, NO SOLUTION

OpenStudy (anonymous):

But u need to PROVE it

OpenStudy (anonymous):

We sure? Do you have a explanation/proof to that?

OpenStudy (anonymous):

Ahh okay

OpenStudy (anonymous):

I guess I have got one proof...

OpenStudy (shubhamsrg):

BY no solution, you mean no integral solution right ?

OpenStudy (anonymous):

The proof that I have got is very simple

OpenStudy (anonymous):

Shoot.

OpenStudy (anonymous):

I will message u if u want

OpenStudy (anonymous):

I'd like to see it. I feel there is a solution as n^n^n grows faster than a quadratic and the quadratic is smaller than the exponent around n=0

OpenStudy (shubhamsrg):

WOlfram tells me a solution, You must be meaning no integral solutions are there hmm ?

OpenStudy (anonymous):

How did you enter that into wolfram? I couldnt get it to work.

OpenStudy (anonymous):

Sorry, I mean integer solution of n

OpenStudy (shubhamsrg):

hmm, that is what i assumed.

OpenStudy (anonymous):

Well from the question n^n +n^(n-1) + n^(n-2)+...+n^2+n^1+n^0 = n^n^n it is clear n>=0 and n is integer

OpenStudy (anonymous):

So yea we agreed there is no solution? as the only intersection is irrational.

OpenStudy (anonymous):

Now, u just need to prove it

OpenStudy (anonymous):

Surely this would suffice: n^n +n^(n-1) + n^(n-2)+...+n^2+n^1+n^0>n^n^n \[\forall n \le 1\] and n^n +n^(n-1) + n^(n-2)+...+n^2+n^1+n^0<n^n^n \[\forall n ge 2\] Q.E.D

OpenStudy (anonymous):

ge is supposed to read as >=

OpenStudy (anonymous):

But u directly said it without any proof

OpenStudy (anonymous):

Let A=n^n +n^(n-1) + n^(n-2)+...+n^2+n^1+n^0, and let B=n^n^n. Note: n is belonging arbitrarily to the set of integers. Then since A<B for all n>1, i.e. all n>=2, and A>B for all n<=1 then since there exists no integer between 1 and 2, there is no possible integer solutions to A=B

OpenStudy (anonymous):

Since R.H.S is divisible by n, L.H.S must also be divisible by n. Now, all the terms except n^0 of L.H.S is divisible by n for all values of n. So,L.H.S is divisible by n only if n^0 is also divisible by n. n^0 is divisible by n only when n=1 but n=1 is not the solution. THUS, NO SOLUTION

OpenStudy (anonymous):

Interesting how you look at it mroe from the abstract whilst i can see it only from the geometric side of things.

OpenStudy (shubhamsrg):

neatly done @sauravshakya